# Auriga Streams III: the mass-metallicity relation does not rule out tidal mass loss in Local Group satellites

### v1

@@ Abstract
Context. The mass-metallicity relation is a fundamental galaxy scaling law that has been extended to the faintest systems in the Local Group. The small scatter in this relation has been used to argue against severe tidal mass loss in Local Group satellites.
Aims. We aim to show that the observed scatter is consistent with tidal mass loss seen in cosmological hydrodynamical zoom-in simulations, using simulated Milky Way-mass haloes from the Auriga project as an example.
Methods. We analyze the stellar masses and mean metallicites for all accreted systems in Auriga. We compute these quantities two ways: considering the total system (bound + lost material) and only considering the bound progenitor. For both cases we fit the mass-metallicity relation and consider properties that drive scatter about these relations.
Results. Accreted systems in Auriga have a tight relation between total stellar mass and metallicity, with scatter at a fixed stellar mass driven by age. When only considering the bound progenitor, the tidally evolved mass-metallicity relation has similar scatter ($\sim$0.27 dex) as observed for the Milky Way and M31 satellites ($\sim$0.23 dex). Satellites that lie above the relation have experienced substantial mass loss and typically have low metallicity for their total stellar mass. Even satellites that fall exactly on the evolved relation can lose over half of their stellar mass. Only satellites that are substantially below the evolved relation are reliably intact.
Conclusions. The small scatter in the observed mass-metallicity relation is compatible with tidal mass loss. Based on their offset from the observed relation, we predict which Milky Way and M31 satellites have experienced substantial mass loss. These predictions will be tested with deep resolved-star observations from upcoming facilities including Rubin LSST, Euclid, and Roman.
Key Words.:
@@ 4 The Local Group satellites
  • Class 3 (0.05 ¡ $f_{\text{bound}}$ ¡ 0.8): Fornax, Leo II, NGC 147, And I, II, III, VI, XIX, XXI, XXVIII, XXIX
@@ 5 Concluding remarks
Acknowledgements.
We thank Astha Chaturvedi, Emma Dodd, Azi Fattahi, Peter Ferguson, and Guilherme Limberg for enlightening discussions, as well as Jess Doppel and Louise Welsh for peer pressure to ‘just write the damn paper’. AHR thanks Taylor Swift’s 1989 (Taylor’s Version) and Bring Me The Horizon’s POST HUMAN: NeX GEn for providing a soundscape condusive to writing this manuscript. We have used simulations from the Auriga Project public data release (Grand et al., 2024) available on the Globus platform [^9] https://wwwmpa.mpa-garching.mpg.de/auriga/data. This research made extensive use of arXiv.org and NASA’s Astrophysics Data System for bibliographic information.
@@ Acknowledgements.
This research made use of the Python programming language, along with many community-developed or maintained software packages including: astropy (Astropy Collaboration et al., 2013, 2018, 2022), cmasher (van der Velden, 2020; van der Velden et al., 2024) cython (Behnel et al., 2011) h5py (Collette, 2013; Collette et al., 2022) jupyter (Perez & Granger, 2007; Kluyver et al., 2016) matplotlib (Hunter, 2007) numpy (Harris et al., 2020) pandas (McKinney, 2010; pandas development team, 2022) scikit-learn (Pedregosa et al., 2011; Buitinck et al., 2013; Grisel et al., 2022) scipy (Virtanen et al., 2020; Gommers et al., 2022). We also thank the maintainers of the arepo-snap-util package. Parts of the results in this work make use of the colormaps in the cmasher package. Software citation information aggregated using The Software Citation Station (Wagg & Broekgaarden, 2024; Wagg et al., 2024).
@@ Appendix B Disruption prediction for observed satellites
$\text{[Fe/H]}=-1.80+0.35\log_{\text{10}}\left(\frac{M_{\star}}{10^{6}\text{\leavevmode\nobreak\ M${}_{\odot}$}}\right)$ (2)
@@
$\text{[Fe/H]}=-1.77+0.29\log_{\text{10}}\left(\frac{M_{\star}}{10^{6}\text{\leavevmode\nobreak\ M${}_{\odot}$}}\right)$ (3)
@@
We note that there are three M31 satellites (M 32, IC 10, and LGS 3) brighter than $M_{V}$ = 7.7-7.7 omitted from our analysis because they lack literature spectroscopic metallicities in the compilations we consider. The available isochrone or RGB colour metallicities compiled by McConnachie (2012) suggest that IC 10 and LGS 3 are not candidates for heavy disruption under our framework (Class 1 or 2 depending on the adopted relation). However, the available data for M 32 gives it a very high metallicity ($\text{[Fe/H]}=-0.25$; Grillmair et al. 1996) for its stellar mass. This would place M 32 in Class 5, comparable to Sagittarius.

### v2

@@ Abstract
The mass-metallicity relation is a fundamental galaxy scaling law that has been extended to the faintest systems in the Local Group. We show that the small scatter in this relation, which has been used to argue against tidal mass loss in Local Group satellites, is consistent with the level of disruption in the Auriga simulations. For every accreted system in Auriga, we compute stellar masses and metallicities two ways: considering the total system (bound + lost material) and only considering the progenitor. Accreted systems in Auriga have a tight relation between total stellar mass and metallicity, with scatter at a fixed stellar mass driven by age. When only considering the progenitor, the tidally evolved mass-metallicity relation has similar scatter ($\sim$0.27 dex) as observed for the Local Group satellites ($\sim$0.23 dex). Satellites that lie above the relation have experienced substantial mass loss and typically have low metallicity for their total stellar mass. Even satellites that fall exactly on the evolved relation can lose over half of their stellar mass. Only satellites substantially below the evolved relation are reliably intact. Based on their offset from the observed relation, we predict which Milky Way and M31 satellites have tidal tails waiting to be discovered.
keywords:
@@
pubyear: 2025pagerange: Auriga Streams III: the mass-metallicity relation does not rule out tidal mass loss in Local Group satellitesB
@@ 4 The Local Group satellites
  • Class 3 (0.05 < $f_{\text{bound}}$ < 0.8): Fornax, Leo II, NGC 147, And I, II, III, VI, XIX, XXI, XXVIII, XXIX
@@ 5 Concluding remarks
Acknowledgements
We thank Astha Chaturvedi, Emma Dodd, Azi Fattahi, Peter Ferguson, and Guilherme Limberg for enlightening discussions, as well as Jess Doppel and Louise Welsh for peer pressure to ‘just write the damn paper’. AHR thanks Taylor Swift’s 1989 (Taylor’s Version) and Bring Me The Horizon’s POST HUMAN: NeX GEn for providing a soundscape condusive to writing this manuscript. This research made extensive use of arXiv.org and NASA’s Astrophysics Data System for bibliographic information.
@@ Acknowledgements
Software
This research made use of the Python programming language, along with many community-developed or maintained software packages including:
  • (i)
    astropy (Astropy Collaboration et al., 2013, 2018, 2022),
  • (ii)
    cmasher (van der Velden, 2020; van der Velden et al., 2024)
  • (iii)
    cython (Behnel et al., 2011)
  • (iv)
    h5py (Collette, 2013; Collette et al., 2022)
  • (v)
    jupyter (Perez & Granger, 2007; Kluyver et al., 2016)
  • (vi)
    matplotlib (Hunter, 2007)
  • (vii)
    numpy (Harris et al., 2020)
  • (viii)
    pandas (McKinney, 2010; pandas development team, 2022)
  • (ix)
    scikit-learn (Pedregosa et al., 2011; Buitinck et al., 2013; Grisel et al., 2022)
  • (x)
    scipy (Virtanen et al., 2020; Gommers et al., 2022).
We also thank the maintainers of the arepo-snap-util package. Parts of the results in this work make use of the colormaps in the cmasher package. Software citation information aggregated using The Software Citation Station (Wagg & Broekgaarden, 2024; Wagg et al., 2024).
Data Availability
Halo catalogs, merger trees, and particle data (Section 2) for Auriga are publicly available (detailed in the Auriga project data release; Grand et al., 2024) to download via the Globus platform [^9] https://wwwmpa.mpa-garching.mpg.de/auriga/data. The observational data used in this work is compiled in Table 2 largely thanks to the heroic efforts of Andrew Pace and Alan McConnachie (McConnachie, 2012; Pace, 2024).
@@ Appendix B Disruption prediction for observed satellites
$\text{[Fe/H]}=-1.80+0.35\log_{\text{10}}\left(\frac{M_{\star}}{10^{6}\text{~M${}_{\odot}$}}\right)$ (2)
@@
$\text{[Fe/H]}=-1.77+0.29\log_{\text{10}}\left(\frac{M_{\star}}{10^{6}\text{~M${}_{\odot}$}}\right)$ (3)
@@
We note that there are three M31 satellites (M 32, IC 10, and LGS 3) brighter than $M_{V}$ = 7.7-7.7 omitted from our analysis because they lack literature spectroscopic metallicities in the compilations we consider. The available isochrone or RGB colour metallicities compiled by McConnachie (2012) suggest that IC 10 and LGS 3 are not candidates for heavy disruption under our framework (Class 1 or 2 depending on the adopted relation). However, the available data for M 32 gives it a very high metallicity ($\text{[Fe/H]}=-0.25$; Grillmair et al., 1996) for its stellar mass. This would place M 32 in Class 5, comparable to Sagittarius.